Find the value of x in:
2^(2x) -2^(x+1) -15 = 0

Answer by user_dude2008(1861)

2^(2x) -2^(x+1) -15 = 0


(2^x)^2 -2*2^x -15 = 0

z^2-2z-15=0

(z-5)(z+3)=0

z-5=0 or z+3=0

z=5 or z=-3

2^x=5 or 2^x=-3

2^x=-3 ----> no solution

2^x=5 ---->

Answer: